OpenStudy (anonymous):

implicit differentiation dy/dx of(cos πx + sin πy)4 = 67

zepdrix (zepdrix):

Is the 4 suppose to be an exponent, like this? \[\huge (\cos \pi x+\sin \pi y)^4=67\]

OpenStudy (anonymous):

use the chain rule :)

OpenStudy (abb0t):

1. chain rule: \[\frac{ d }{ dx } a^n = na^{n-1}\] 2. differentiate with respect to x: \[\frac{ dy }{ dx }\] 3. use algebra to rearrange and solve for dy/dx Your final answer should be in terms of only x and y.

OpenStudy (zehanz):

Using the tips from @zepdrix and @abb0t: Differentiate both sides:\[4(\cos \pi x + \sin \pi y)^3(-\pi \sin \pi x + \pi \cos \pi y \frac{ dy }{ dx })=0\]If you divide by 4pi you get:\[(\cos \pi x + \sin \pi y)^3(\frac{ dy }{ dx } \cos \pi y - \sin \pi x)=0\]A product is zero if one or more factors are zero:\[\cos \pi x + \sin \pi y = 0\]or\[\frac{ dy }{ dx }(\cos \pi y - \sin \pi x)=0\]This last one leads to:\[\frac{ dy }{ dx }=\frac{ 1 }{ \cos \pi y - \sin \pi x }\]